# Simultaneous Task Allocation and Planning for Multi-Robots under Hierarchical Temporal Logic Specifications

### v3

@@ VI-A On-the-fly Search
The on-the-fly search algorithm is built upon the Dijkstra’s algorithm, as detailed in Alg. 2. The search state is defined as $v=(r,\phi,{\mathbf{s}},{\mathbf{q}})$, where ${\mathbf{s}}\in{\mathcal{S}}_{1}\otimes\ldots\otimes{\mathcal{S}}_{N}$ and ${\mathbf{q}}={\mathcal{Q}}_{\mathcal{B}}(\phi_{K}^{1})\otimes\ldots\otimes{\mathcal{Q}}_{\mathcal{B}}(\phi_{1}^{1})$. This state is comprised of three parts: (i) $(r,\phi)$ indicates the specific PBA currently being searched; (ii) ${\mathbf{s}}$ represents the states of all robots; and (iii) ${\mathbf{q}}$ denotes the automaton states of all specifications (including leaf and non-leaf ones). Two states are considered equivalent if they share the same $r$, $\phi$, ${\mathbf{s}}$, and ${\mathbf{q}}$ values. The notation ${\mathbf{s}}(r)$ and ${\mathbf{q}}(\phi_{k}^{i})$ refer to the state of robot $r$ and the automaton state of specification $\phi_{k}^{i}$ reached so far, respectively. We use ${\mathcal{V}}_{\text{explored}}$ and ${\mathcal{V}}_{\text{seen}}$ to denote the sets of states that have been explored and seen, and ${\mathcal{H}}_{\text{front}}$ for the states awaiting exploration. States within ${\mathcal{H}}_{\text{front}}$ are ordered by their costs. Lastly, $\Lambda_{\text{explored}}$ maps each explored state to the shortest path that led to it.
@@
  • $(r,\phi,{\mathbf{s}}(r),{\mathbf{q}}(\phi))\in{\mathcal{Q}}_{\mathcal{P}}^{0},\forall r\in[N],\forall\phi\in\Phi^{K}$;
  • ${\mathbf{q}}(\phi)\in{\mathcal{Q}}_{\mathcal{B}}^{0}(\phi),\forall\phi\in\Phi\setminus\Phi^{K}$.
That is, all robots begin at their initial states, and every specification starts from its initial automaton state [line 2]. Whenever a state $v$ is popped out from the set ${\mathcal{H}}_{\text{front}}$ for exploration, it is added to the set of explored states if it hasn’t been already [lines 2-2]. The iteration process concludes when an accepting state of the topmost-level specification is reached. At this point, the function ExtractPlan is employed to parallelize the sequentially searched path, generating a state-specification plan [line 2]. If the accepting state is not yet reached, the algorithm proceeds to determine the successor states of $v$, along with their associated transition costs [line 2], through the function GetSucc, outlined in Alg. 3. For each successor state that is either unseen or has a total cost lower than previously found, it is queued into ${\mathcal{H}}_{\text{front}}$ for future exploration. Concurrently, the corresponding path leading to this state is recorded [lines 2-2].
Input: Hierarchical sc-LTL specifications $\Phi$, a set of robots $[N]$
Output: state-specification plan
@@
2 ${\mathcal{V}}_{\text{init}}=\{v\,|\,(r,\phi,{\mathbf{s}}(r),{\mathbf{q}}(\phi))\in{\mathcal{Q}}_{\mathcal{P}}^{0},\forall r\in[N],\forall\phi\in\Phi^{K}\wedge{\mathbf{q}}(\phi)\in{\mathcal{Q}}_{\mathcal{B}}^{0}(\phi),\forall\phi\in\Phi\setminus\Phi^{K}\}$ ;
@@ 4 $\tau_{\text{explored}}(v)=v,\forall v\in{\mathcal{V}}_{\text{init}}$;
5 Build hierarchy specification graph ${\mathcal{G}}_{h}$;
@@
12       if $v.{\mathbf{q}}(1,1)\in{\mathcal{Q}}_{\mathcal{B}}^{F}(1,1)$ then
@@ 21                   $\Lambda_{\text{explored}}(u)=\Lambda_{\text{explored}}(v)+u$ ;
Input: state $v$, hierarchical team models ${\mathcal{P}}$, hierarchy specification graph ${\mathcal{G}}_{h}$
@@ 1 ${\mathcal{V}}_{\text{succ}}=\varnothing$;
5 if $q\not\in{\mathcal{Q}}^{F}_{\mathcal{B}}(\phi)$ and $v.{\mathbf{q}}(\phi_{\text{prt}})\not\in{\mathcal{Q}}_{\mathcal{B}}^{F}(\phi_{\text{prt}}),\forall\phi_{\text{prt}}\in\Phi_{\text{prt}}$ then
@@ 7             $u=v,{\mathcal{V}}_{u}=\varnothing$;
9             $\pi=\{\phi\}$ if $q^{\prime}\in{\mathcal{Q}}^{F}_{\mathcal{B}}(\phi)$ else $\varnothing$ ;
@@ 14             $u=v$;
18 for $\phi^{\prime}\in\Phi^{K}\setminus\{\phi\}$  do
@@ 20             $u=v$;
24 for $\phi^{\prime}\in\Phi^{K}\setminus\{\phi\}$  do
@@ 26             $u=v$;
The function GetSucc determines the successor states of a given state $v=(r,\phi,{\mathbf{s}},{\mathbf{q}})$ by considering in-spec transitions in ${\mathcal{P}}(r,\phi)$, in-spec switch transitions in ${\mathcal{P}}(v.\phi)$, and inter-spec type I and II switch transitions in ${\mathcal{P}}$. To identify in-spec transitions in ${\mathcal{P}}(r,\phi)$, the function initially checks if the specification $\phi$ or any of its non-leaf parent specifications are satisfied. This is achieved using Parents(${\mathcal{G}}_{h}$, $\phi$), which returns the path from the leaf specification $\phi$ to the topmost specification in the hierarchical specification graph ${\mathcal{G}}_{h}$, excluding $\phi$ itself [line 3]. If $\phi$ or any of its parent specifications are satisfied, the search in ${\mathcal{P}}(r,\phi)$ and ${\mathcal{P}}(\phi)$ ceases, as further exploration is unnecessary. This is based on condition (2) in Definition IV.1 and Assumption IV.6, which state that each leaf specification appears only in one non-leaf specification and satisfying each specification at most once is adequate. If a parent specification is satisfied, the leaf specification no longer contributes to task progress.
Input: True composite propositions $\pi$, state $u$, parent specifications $\Phi_{\text{prt}}$
@@
2 ${\mathcal{Q}}_{\mathcal{B}}^{\text{succ}}=\left\{q^{\prime}\,|\,(u.{\mathbf{q}}(\phi_{\text{prt}}),\pi,q^{\prime})\in\to_{\mathcal{B}}(\phi_{\text{prt}}),\forall q^{\prime}\in{\mathcal{Q}}_{\mathcal{B}}(\phi_{\text{prt}})\right\}$;
3 for $q^{\prime}\in{\mathcal{Q}}^{\text{succ}}_{\mathcal{B}}$ do
@@
9             $\pi=\{\phi_{\text{prt}}\}$ if $q^{\prime}\in{\mathcal{Q}}^{F}_{\mathcal{B}}(\phi_{\text{prt}})$ else $\varnothing$;
@@
If none of the parent specifications are satisfied, the state of robot $r$ in $v.{\mathbf{s}}$ and the automaton state of leaf specification $\phi$ in $v.{\mathbf{q}}$ are updated based on in-spec transitions in $\to_{{\mathcal{P}}(r,\phi)}$ [line 3]. Subsequently, the automaton states of all parent specifications of $\phi$ are also updated, as detailed in lines 3-3 and further elaborated in Alg. 4. For the three types of switch transitions, only the robot $r$ (in the case of in-spec switch transitions [line 3)], the leaf specification $\phi$ (for inter-spec type I switch transitions [line 3]), or both robot $r$ and specification $\phi$ (for inter-spec type II switch transitions [line 3]) are updated. This implies a shift in the search space to a different PBA, without modifying $v.{\mathbf{s}}$ and $v.{\mathbf{q}}$. Since these transitions do not change robot states and automaton states, the function UpdateNonLeafSpecs is not invoked.
Given a state $u$ and the set of parent specifications $\Phi_{\text{prt}}$ of the leaf specification $u.\phi$ currently being explored, the function UpdateNonLeafSpecs utilizes a depth-first strategy to refresh the automaton states of non-leaf specifications within $\Phi_{\text{prt}}$. When considering a specific parent specification $\phi_{\text{prt}}\in\Phi_{\text{prt}}$, the function initially identifies all states that can be reached based on the acceptance of the preceding specification at the immediate lower level [lines 4-4]. The automaton state of $\phi_{\text{prt}}$ in $u.{\mathbf{q}}$ is then updated using these reachable states [line 4]. This process continues until the automaton state of the topmost specification is updated [line 4].
@@
Output: state-specification plan $\tau$
@@
4             $\tau^{i}=((s_{1}^{i},\phi_{1}^{i}),\ldots,(s_{r}^{i},\phi_{r}^{i}),\ldots,(s_{N}^{i},\phi_{N}^{i}))$ where $(s_{r}^{i},\phi_{r}^{i})=(v.{\mathbf{s}}(r),v.\phi)$ and $(s_{r^{\prime}}^{i},\phi_{r^{\prime}}^{i})=(v.{\mathbf{s}}(r^{\prime}),\epsilon),\forall r^{\prime}\not=v.r$ ;
@@
The ExtractPlan function in Alg. 5 processes the path $\Lambda=v_{0}v_{1}\ldots v_{n}$, which leads to the state satisfying the topmost specification, and outputs a state-specification plan $\tau$. This process is visually represented in Fig. 4. A state $v$ within $\Lambda$ is identified as a switch state if it is the end state of an inter-spec switch transition $\zeta_{\text{inter}}^{1}\cup\zeta_{\text{inter}}^{2}$, which has the same robot state and automaton state as the corresponding start state. The path $\Lambda$ is segmented by these switch states, with each segment corresponding to states associated with the same specification, indicative of a search within a specific product team model. To create a concise path $\Lambda^{\prime}$, all switch states are removed, as they merely facilitate transitions between product team models and don’t influence the plan’s progress. This refined path is denoted as $\Lambda^{\prime}=\Lambda_{1}\Lambda_{2}\ldots\Lambda_{l}$, where $\Lambda_{k}$ represents the $k$-th path segment, as shown in line 5. The next step involves converting these sequential path segments into parallel executions. For each segment $\Lambda_{k}$, we identify all active robots ${\mathcal{R}}_{k}$ participating in the specification $\phi_{k}$, represented as ${\mathcal{R}}_{k}=\left\{v.r\,|\,v\in\Lambda_{k}\right\}$. Due to task decomposition [8], each robot in ${\mathcal{R}}_{k}$ can execute its part of the plan independently. Robots not in ${\mathcal{R}}_{k}$ are considered inactive for that segment. The plan horizon post-parallelization is determined by the longest path among all active robots. Note that an active robot might not maintain the active mode during the entire horizon if it completes its actions in a shorter time frame. For each robot state, we pair it with the relevant specification $\phi_{k}$ if the robot is active, or with a null specification $\epsilon$ if it’s inactive [line 5].
@@
Our method aligns with the decoupled approaches in the domain MAPF [68, 69], where initially, individual agents independently plan their paths without considering potential collisions. Subsequently, the system assesses for collisions and modifies the paths to prevent conflicts. Although this approach is more scalable and demands less computational effort than fully coupled methods, it may not ensure an optimal outcome, could lead to deadlocks, and might face challenges in dense areas. Resolving conflicts in temporal logic planning proves more difficult because, unlike MAPF, the actions of robots may be temporally correlated. When addressing collisions, it is crucial not only to consider robots that are physically close but also those that are temporally correlated. Therefore, we leave the systemic exploration of a complete collision resolution strategy to future studies.
@@ VI-B2 Essential Switch Transitions
We introduce essential switch transitions to refine the connections between PBAs.
Definition VI.3 (Essential State)
@@
  • $q^{\prime}\in{\mathcal{D}}_{\mathcal{B}}(\phi)$.
@@
Definition VI.4 (Essential Switch Transition)
@@ VI-B3 Guidance by Automaton States
$c_{q}({\mathbf{q}})&=\sum_{\phi^{\prime}\in\Phi^{K}}c_{q}({\mathbf{q}}(\phi^{\prime}))\\ =\sum_{\phi^{\prime}\in\Phi^{K}}&\max\left\{\texttt{len}(q^{0}_{\mathcal{B}}\to{\mathbf{q}}(\phi^{\prime}))\,|\,\forall\,q^{0}_{\mathcal{B}}\in{\mathcal{Q}}_{\mathcal{B}}^{0}(\phi^{\prime})\right\}.\end{aligned}$ (4)
The cost $c_{q}({\mathbf{q}}(\phi^{\prime}))$ is determined by the length of the simplest path from any initial state $q^{0}_{\mathcal{B}}$ to the automaton state ${\mathbf{q}}(\phi^{\prime})$ within the NBA ${\mathcal{B}}(\phi)$. A simple path is defined as one that does not include any repeated states, and the function len measures the path’s length in terms of the number of transitions. Essentially, $c_{q}({\mathbf{q}}(\phi))$ captures the maximal progress made in the task $\phi$ since its beginning. A similar heuristic has been employed in [18, 21]. The cost $c_{vu}$ is now calculated as
@@
Remark VI.5
The first two heuristics essentially alter the structure of the hierarchical team models. As a result, the completeness and optimality of the search approach, when these heuristics are applied, are no longer assured. However, employing exclusively the third heuristic still maintains the completeness of the search, as it ensures the full exploration of the search space eventually. Additionally, this approach with heuristic of automaton state remains optimal with respect to the newly defined cost (5). Extensive simulations presented in Sec. VIII provide empirical evidence that our approach, incorporating all three heuristics, can generate solutions for complex tasks. The rationale behind this empirical completeness is twofold. Firstly, [28] prove that the inference on temporal order is correct for a broad range of flat sc-LTL specifications. By eliminating switch transitions that conflict with precedence orders, transitions that are in accordance with temporal orders remain unaffected. Secondly, the concept of essential switch transitions ensures that the search process only transitions to a different product model when there has been task progress in the current product model, which is reasonable as it is unnecessary to switch to a different task if the ongoing task has not been completed.

### v4

@@ VI-A On-the-fly Search
The on-the-fly search algorithm is built upon the Dijkstra’s algorithm, as detailed in Alg. 2. The search state is defined as $v=(r,\phi,{\mathbf{s}},{\mathbf{q}})$, where ${\mathbf{s}}\in{\mathcal{S}}_{1}\otimes\ldots\otimes{\mathcal{S}}_{N}$ and ${\mathbf{q}}={\mathcal{Q}}_{\mathcal{A}}(\phi_{K}^{1})\otimes\ldots\otimes{\mathcal{Q}}_{\mathcal{A}}(\phi_{1}^{1})$. This state is comprised of three parts: (i) $(r,\phi)$ indicates the specific PA currently being searched; (ii) ${\mathbf{s}}$ represents the states of all robots; and (iii) ${\mathbf{q}}$ denotes the automaton states of all specifications (including leaf and non-leaf ones). Two states are considered equivalent if they share the same $r$, $\phi$, ${\mathbf{s}}$, and ${\mathbf{q}}$ values. The notation ${\mathbf{s}}(r)$ and ${\mathbf{q}}(\phi_{k}^{i})$ refer to the state of robot $r$ and the automaton state of specification $\phi_{k}^{i}$ reached so far, respectively. We use ${\mathcal{V}}_{\text{explored}}$ and ${\mathcal{V}}_{\text{seen}}$ to denote the sets of states that have been explored and seen, and ${\mathcal{H}}_{\text{front}}$ for the states awaiting exploration. States within ${\mathcal{H}}_{\text{front}}$ are ordered by their costs. Lastly, $\Lambda_{\text{explored}}$ maps each explored state to the shortest path that led to it.
@@
  • $(r,\phi,{\mathbf{s}}(r),{\mathbf{q}}(\phi))\in{\mathcal{Q}}_{\mathcal{P}}^{0},\forall r\in[N]_{+},\forall\phi\in\Phi_{\text{leaf}}$;
  • ${\mathbf{q}}(\phi)\in{\mathcal{Q}}_{\mathcal{A}}^{0}(\phi),\forall\phi\in\Phi\setminus\Phi_{\text{leaf}}$.
That is, all robots begin at their initial states, and every specification starts from its initial automaton state [line 2]. Whenever a state $v$ is popped out from the set ${\mathcal{H}}_{\text{front}}$ for exploration, it is added to the set of explored states if it hasn’t been already [lines 2-2]. The iteration process concludes when an accepting state of the topmost-level specification is reached. At this point, the function ExtractPlan is employed to parallelize the sequentially searched path, generating a state-specification sequence [line 2]. If the accepting state is not yet reached, the algorithm proceeds to determine the successor states of $v$, along with their associated transition costs [line 2], through the function GetSucc, outlined in Alg. 3. For each successor state that is either unseen or has a total cost lower than previously found, it is queued into ${\mathcal{H}}_{\text{front}}$ for future exploration. Concurrently, the corresponding path leading to this state is recorded [lines 2-2].
Input: H-LTLf specifications $\Phi$, a set of robots $[N]_{+}$
Output: state-specification sequence
@@
2 ${\mathcal{V}}_{\text{init}}=\{v\,|\,(r,\phi,{\mathbf{s}}(r),{\mathbf{q}}(\phi))\in{\mathcal{Q}}_{\mathcal{P}}^{0},\forall r\in[N]_{+},\forall\phi\in\Phi_{\text{leaf}}\wedge{\mathbf{q}}(\phi)\in{\mathcal{Q}}_{\mathcal{A}}^{0}(\phi),\forall\phi\in\Phi\setminus\Phi_{\text{leaf}}\}$ ;
@@ 4 $\tau_{\text{explored}}(v)=v,\forall v\in{\mathcal{V}}_{\text{init}}$;
5 Build specification hierarchy tree ${\mathcal{G}}_{h}$;
@@
12       if $v.{\mathbf{q}}(\phi_{1}^{1})\in{\mathcal{Q}}_{\mathcal{A}}^{F}(\phi_{1}^{1})$  then
@@ 21                   $\Lambda_{\text{explored}}(u)=\Lambda_{\text{explored}}(v)+u$ ;
Input: state $v$, hierarchical team models ${\mathcal{P}}$, specification hierarchy tree ${\mathcal{G}}_{h}$
@@ 1 ${\mathcal{V}}_{\text{succ}}=\varnothing$;
5 if $q\not\in{\mathcal{Q}}^{F}_{\mathcal{A}}(\phi)$ and $v.{\mathbf{q}}(\phi_{\text{prt}})\not\in{\mathcal{Q}}_{\mathcal{A}}^{F}(\phi_{\text{prt}}),\forall\phi_{\text{prt}}\in\Phi_{\text{prt}}$ then
@@ 7             $u=v,{\mathcal{V}}_{u}=\varnothing$;
9             $\pi_{c}=\{\phi\}$ if $q^{\prime}\in{\mathcal{Q}}^{F}_{\mathcal{A}}(\phi)$ else $\varnothing$ ;
@@ 14             $u=v$;
18 for $\phi^{\prime}\in\Phi_{\text{leaf}}\setminus\{\phi\}$  do
@@ 20             $u=v$;
24 for $\phi^{\prime}\in\Phi_{\text{leaf}}\setminus\{\phi\}$  do
@@ 26             $u=v$;
The function GetSucc determines the successor states of a given state $v=(r,\phi,{\mathbf{s}},{\mathbf{q}})$ by considering in-spec transitions in ${\mathcal{P}}(r,\phi)$, in-spec switch transitions in ${\mathcal{P}}(v.\phi)$, and inter-spec type I and II switch transitions in ${\mathcal{P}}$. To identify in-spec transitions in ${\mathcal{P}}(r,\phi)$, the function initially checks if the specification $\phi$ or any of its non-leaf parent specifications are satisfied. This is achieved using Parents(${\mathcal{G}}_{h}$, $\phi$), which returns the path from the leaf specification $\phi$ to the root specification in the specification hierarchy tree ${\mathcal{G}}_{h}$, excluding $\phi$ itself [line 3]. If $\phi$ or any of its parent specifications are satisfied, the search in ${\mathcal{P}}(r,\phi)$ and ${\mathcal{P}}(\phi)$ ceases, as further exploration is unnecessary. This is based on condition (3) in Definition IV.1 and Assumption V.1, which state that each leaf specification appears only in one non-leaf specification and satisfying each specification at most once is adequate.
Input: True composite propositions $\pi_{c}$, state $u$, parent specifications $\Phi_{\text{prt}}$
@@
2 ${\mathcal{Q}}_{\mathcal{A}}^{\text{succ}}=\left\{q^{\prime}\,|\,(u.{\mathbf{q}}(\phi_{\text{prt}}),\pi_{c},q^{\prime})\in\to_{\mathcal{A}}(\phi_{\text{prt}}),\forall q^{\prime}\in{\mathcal{Q}}_{\mathcal{A}}(\phi_{\text{prt}})\right\}$;
3 for $q^{\prime}\in{\mathcal{Q}}^{\text{succ}}_{\mathcal{A}}$ do
@@
9             $\pi_{c}=\{\phi_{\text{prt}}\}$ if $q^{\prime}\in{\mathcal{Q}}^{F}_{\mathcal{A}}(\phi_{\text{prt}})$ else $\varnothing$;
@@
If none of the parent specifications are satisfied, the state of robot $r$ in $v.{\mathbf{s}}$ and the automaton state of leaf specification $\phi$ in $v.{\mathbf{q}}$ are updated based on in-spec transitions in $\to_{{\mathcal{P}}(r,\phi)}$ [line 3]. Subsequently, the automaton states of all parent specifications of $\phi$ are also updated, as detailed in lines 3-3 and further elaborated in Alg. 4. For the three types of switch transitions, only the robot $r$ (in the case of in-spec switch transitions [line 3)], the leaf specification $\phi$ (for inter-spec type I switch transitions [line 3]), or both robot $r$ and specification $\phi$ (for inter-spec type II switch transitions [line 3]) are updated. This implies a shift in the search space to a different PA, without modifying $v.{\mathbf{s}}$ and $v.{\mathbf{q}}$. Since these transitions do not change robot states and automaton states, the function UpdateNonLeafSpecs is not invoked.
Given a state $u$ and the set of parent specifications $\Phi_{\text{prt}}$ of the leaf specification $u.\phi$ currently being explored, the function UpdateNonLeafSpecs utilizes a depth-first strategy to refresh the automaton states of non-leaf specifications within $\Phi_{\text{prt}}$. When considering a specific parent specification $\phi_{\text{prt}}\in\Phi_{\text{prt}}$, the function initially identifies all states that can be reached based on the acceptance of the preceding specification at the immediate lower level [lines 4-4]. The automaton state of $\phi_{\text{prt}}$ in $u.{\mathbf{q}}$ is then updated using these reachable states [line 4]. This process continues until the automaton state of the root specification is updated [line 4].
@@
Output: state-specification sequence $\tau$
@@
4             $\tau_{i}=((s_{1}^{i},\psi_{1}^{i}),\ldots,(s_{r}^{i},\psi_{r}^{i}),\ldots,(s_{N}^{i},\psi_{N}^{i}))$ where $(s_{r}^{i},\psi_{r}^{i})=(v.{\mathbf{s}}(r),v.\phi)$ and $(s_{r^{\prime}}^{i},\psi_{r^{\prime}}^{i})=(v.{\mathbf{s}}(r^{\prime}),\epsilon),\forall r^{\prime}\not=v.r$ ;
@@
The ExtractPlan function in Alg. 5 processes the path $\Lambda=v_{0}v_{1}\ldots v_{n}$, which leads to the state satisfying the root specification, and outputs a state-specification sequence $\tau$. This process is visually represented in Fig. 5. A state $v$ within $\Lambda$ is identified as a switch state if it is the end state of an inter-spec switch transition $\zeta_{\text{inter}}^{1}\cup\zeta_{\text{inter}}^{2}$, which has the same robot state and automaton state as the corresponding start state. The path $\Lambda$ is segmented by these switch states, with each segment corresponding to states associated with the same specification, indicative of a search within a specific product team model. To create a concise path $\Lambda^{\prime}$, all switch states are removed, as they merely facilitate transitions between product team models and don’t influence the plan’s progress. This refined path is denoted as $\Lambda^{\prime}=\Lambda_{1}\Lambda_{2}\ldots\Lambda_{l}$, where $\Lambda_{k}$ represents the $k$-th path segment, as shown in line 5. The next step involves converting these sequential path segments into parallel executions. For each segment $\Lambda_{k}$, we identify all active robots ${\mathcal{R}}_{k}$ participating in the specification $\phi_{k}$, represented as ${\mathcal{R}}_{k}=\left\{v.r\,|\,v\in\Lambda_{k}\right\}$. Due to task decomposition [9], each robot in ${\mathcal{R}}_{k}$ can execute its part of the plan independently. Robots not in ${\mathcal{R}}_{k}$ are considered inactive for that segment. The plan horizon post-parallelization is determined by the longest path among all active robots. Note that an active robot might not maintain the active mode during the entire horizon if it completes its actions in a shorter time frame. For each robot state, we pair it with the relevant specification $\phi_{k}$ if the robot is active, or with a null specification $\epsilon$ if it’s inactive [line 5].
@@
The search state includes the joint robot states to keep track of the states of idle robots, allowing the search to resume from their previous states when necessary. In Fig. 4, after completing the search for robot 2 with respect to specification $\phi_{1}$, the search transitions to robot 1 while robot 2 remains idle. It then returns to robot 2 for specification $\phi_{2}$ via an in-spec switch transition ③. At this stage, knowing the state of robot 2 is essential.
Remark VI.3
Our method aligns with the decoupled approaches in the domain MAPF [67, 68], where initially, individual agents independently plan their paths without considering potential collisions. Subsequently, the system assesses for collisions and modifies the paths to prevent conflicts. Resolving conflicts in temporal logic planning proves more difficult because, unlike MAPF, the actions of robots may be temporally correlated. When addressing collisions, it is crucial not only to consider robots that are physically close but also those that are temporally correlated. Moreover, the modified path must be carefully adjusted to ensure that the temporal logic task remains satisfied. Therefore, we leave the exploration of a collision resolution strategy to future studies.
@@ VI-B2 Essential Switch Transitions
We introduce essential switch transitions to refine the connections between PAs.
Definition VI.4 (Essential State)
@@
  • $q^{\prime}\in{\mathcal{D}}_{\mathcal{A}}(\phi)$.
@@
Definition VI.5 (Essential Switch Transition)
@@ VI-B3 Guidance by Automaton States
$c_{q}({\mathbf{q}})&=\sum_{\phi^{\prime}\in\Phi_{\text{leaf}}}c_{q}({\mathbf{q}}(\phi^{\prime}))\\ =\sum_{\phi^{\prime}\in\Phi_{\text{leaf}}}&\max\left\{\texttt{len}(q^{0}_{\mathcal{A}}\to{\mathbf{q}}(\phi^{\prime}))\,|\,\forall\,q^{0}_{\mathcal{A}}\in{\mathcal{Q}}_{\mathcal{A}}^{0}(\phi^{\prime})\right\}.\end{aligned}$ (5)
The cost $c_{q}({\mathbf{q}}(\phi^{\prime}))$ is determined by the length of the simplest path from any initial state $q^{0}_{\mathcal{A}}$ to the automaton state ${\mathbf{q}}(\phi^{\prime})$ within the NFA ${\mathcal{A}}(\phi)$. A simple path is defined as one that does not include any repeated states, and the function len measures the path’s length in terms of the number of transitions. Essentially, $c_{q}({\mathbf{q}}(\phi))$ captures the maximal progress made in the task $\phi$ since its beginning. A similar heuristic has been employed in [19, 22]. The cost $c_{vu}$ is now calculated as
@@
Remark VI.6
The first two heuristics essentially alter the structure of the hierarchical team models. As a result, the completeness and optimality of the search approach, when these heuristics are applied, are no longer assured. However, employing exclusively the third heuristic still maintains the completeness of the search, as it ensures the full exploration of the search space eventually. Extensive simulations presented in Sec. VIII provide empirical evidence that our approach, incorporating all three heuristics, can generate solutions for complex tasks. The rationale behind this empirical completeness is twofold. Firstly, [29] prove that the inference on temporal order is correct for a broad range of standard LTL specifications. By eliminating switch transitions that conflict with precedence orders, transitions that are in accordance with temporal orders remain unaffected. Secondly, the concept of essential switch transitions ensures that the search process only transitions to a different product model when there has been task progress in the current product model, which is reasonable as it is unnecessary to switch to a different task if the ongoing task has not been completed.